OpenStudy (anonymous):

Please oh please help me with this one! Graph the feasible set for the system of inequalities x+2y>= 6 x+y>=5 x>=1

OpenStudy (anonymous):

?

OpenStudy (triciaal):

this should help eqn 3 x>=1 draw the line and shade to the right eqn 2 when x = 1 y= 4 y = 5 - x find the section y>5-x; x = 2 y = 3 ;x = 0 y=5 when y = 0 x = 5 eqn 1 when x = 0 y =6/2=3 ; x = 1 y = (6-1)/2 =5/2 when y = 0 x =6 plot the points draw the lines and find the overlapping section.

OpenStudy (anonymous):

For each equation solve for y and x by using x=0 then solve for y =0 then use substituttion then you will get x and y coordinate for that equation. do each equation this way and draw a line for each one. then for feasibility you test the areas that are not on the line to see if that number holds true for each equation.